Suppose V is finite-dimensional with dim V greater or equal to 2. Prove that

there exist S; T € L(V; V ) such that ST is not equal to T S.


1
Expert's answer
2021-07-14T08:39:10-0400

Let "v_1,v_2,...,v_n" be a basis of "V".

Define "S,T\in L(V,V)" such that

"S(v_l)=
\begin{cases} 
 v_2 & \text{if }l=1 \\
 0 & otherwise
 \end{cases} \\
T(v_l)=
\begin{cases} 
 v_1 & \text{if }l=2 \\
 0 & otherwise
 \end{cases}"

Consider

"(ST)(v_1)=S(T(v_1))=S(0)=0", since "S" is linear.

But

"(TS)(v_1)=T(S(v_1))=T(v_2)=v_1 \neq 0"

So "ST \neq TS".